OpenStudy (anonymous):

if a die is rolled twice, what is probability of rolling a 5 and then a 2? @misty1212 @misty1212 @misty1212

OpenStudy (anonymous):

If there is one face with 5 and one face with 2 on a die, and there are 6 faces total, then the probability for each will be 1/6 When you roll a die, it does not affect the next roll, so it's going to be "the probability of this AND this" To find this type, you multiply your ratios. Do you think you can solve it from here?

OpenStudy (anonymous):

no im stupid >-<

OpenStudy (anonymous):

That's all right. Your ratios are 1/6 and 1/6, so your multiplication will look like this \[\frac{ 1 }{ 6 }\times\frac{ 1 }{ 6 }\] Then you cross multiply (NumeratorxNumerator, DenominatorxDenominator)

OpenStudy (anonymous):

Does that make sense/do you think you can solve it from here?

OpenStudy (anonymous):

im not good with math at all >~<

OpenStudy (anonymous):

nm is 1/36

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

Yeah! =D

OpenStudy (anonymous):

can it be simplified

OpenStudy (anonymous):

No, because 1 is not divisible by anything.

OpenStudy (anonymous):

will you help me somemore

OpenStudy (anonymous):

Sure =)